# Federated Distributed Key Generation

### v2

@@ III-A Notation
Let $\lambda$ denote the computational security parameter. We work within a cyclic group $\mathbb{G}$ of prime order $q$ generated by $G\in\mathbb{G}$. Group operations are written multiplicatively, and scalar multiplication is denoted by $G^{a}$ for $a\in\mathbb{Z}_{q}$. We assume the Discrete Logarithm Problem (DLP) is hard in $\mathbb{G}$. All algorithms are assumed to be Probabilistic Polynomial-Time (PPT) unless stated otherwise. Sampling a random element $x$ from a set $S$ is denoted by $x\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,S$. We denote the set of potential parties as $\mathbb{P}=\{P_{1},\dots,P_{n}\}$.
@@ III-C Verifiable and Publicly Verifiable Secret Sharing (VSS/PVSS)
VSS schemes enhance SSS by allowing parties to verify the consistency of their shares relative to some public information. PVSS strengthens this by enabling any entity to perform this verification using only public data [12, 13]. PVSS schemes typically involve the dealer encrypting each share $s_{i}$ under the recipient $P_{i}$’s public key $\textrm{pk}_{i}$ and publishing these ciphertexts $C_{i}=\texttt{Enc}(\textrm{pk}_{i},s_{i})$ along with public commitments to the secret polynomial (e.g., $\{A_{k}=G^{a_{k}}\}$) and a public proof $\pi$ of consistency. Our FDKG protocol employs a PVSS scheme where each participant acts as a dealer.
@@ III-D Public Key Encryption (PKE)
  • $(\textrm{pk}_{i},\textrm{sk}_{i})\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,\textrm{KeyGen}(1^{\lambda})$: Generates a public key $\textrm{pk}_{i}$ and a secret key $\textrm{sk}_{i}$.
@@ III-E Non-Interactive Zero-Knowledge (NIZK) Proofs
  • $\textrm{crs}\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,\textrm{Setup}(1^{\lambda},\mathcal{R})$: Generates the public CRS. (A simulator might receive additional trapdoor information during security analysis, depending on the scheme).
  • $\pi\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,\textrm{Prove}(\textrm{crs},x,w)$: Produces proof $\pi$.
@@ III-F Encrypted Channels for Share Distribution
Protocols like PVSS and DKG often require confidential transmission of secret shares (e.g., $s_{i}$) to specific recipients ($P_{i}$) over a public broadcast channel. This is achieved using an IND-CPA secure PKE scheme. Each potential recipient $P_{i}$ has a key pair $(\textrm{pk}_{i},\textrm{sk}_{i})$ where $\textrm{pk}_{i}$ is public. To send $s_{i}$ to $P_{i}$, the sender computes $c=\texttt{Enc}(\textrm{pk}_{i},s_{i},\textrm{rand})$ and broadcasts $c$. Only $P_{i}$ can compute $s_{i}=\texttt{Dec}(\textrm{sk}_{i},c)$.
@@
1 Input: Recipient public key $\textrm{pk}_{i}$, scalar message $m\in\mathbb{Z}_{q}$, randomness $(k,r)\in\mathbb{Z}_{q}\times\mathbb{Z}_{q}$
@@ 5 $C_{2}=(\textrm{pk}_{i})^{k}\cdot M$
Algorithm 1 ElGamal Encryption $\texttt{Enc}(\textrm{pk}_{i},m,[k,r])$
1 Input: Recipient secret key $\textrm{sk}_{i}$, ciphertext $(C_{1},C_{2},\Delta)$
@@ 3 $M=C_{2}\cdot(C_{1}^{\textrm{sk}_{i}})^{-1}$
Algorithm 2 ElGamal Decryption $\texttt{Dec}(\textrm{sk}_{i},(C_{1},C_{2},\Delta))$
@@ III-G Summary of Key Notations
Table I: Key Notations Used in the FDKG Protocol
@@
$\textrm{pk}_{i},\textrm{sk}_{i}$ PKE public/secret key pair for $P_{i}$ (for encrypted channels).
@@
$t$ Local threshold for SSS reconstruction within a guardian set.
@@
$\mathbb{D}$ Set of parties participating and validated in the FDKG phase.
@@
$\texttt{Enc}(\textrm{pk},\cdot,\cdot)$ PKE encryption algorithm.
$\texttt{Dec}(\textrm{sk},\cdot)$ PKE decryption algorithm.
@@
$B_{i},v_{i},r_{i}$ Voter $P_{i}$’s encrypted ballot, encoded vote, and blinding factor.
@@ IV Federated Distributed Key Generation (FDKG)
Federated Distributed Key Generation (FDKG) extends traditional Distributed Key Generation (DKG) protocols to accommodate settings with dynamic participation and participant-defined trust assumptions. Standard $(t,n)$-DKG protocols require a fixed set of $n$ participants and a global threshold $t$ for secret reconstruction [2]. FDKG relaxes these requirements, allowing any subset of potential participants to engage in key generation, each delegating reconstruction authority to a self-chosen ”guardian set”.
@@ IV-A Protocol Setting and Goal
We consider a set $\mathbb{P}=\{P_{1},\dots,P_{n}\}$ of $n$ potential parties. The system operates over a cyclic group $\mathbb{G}$ of prime order $q$ with generator $G$. We assume a Public Key Infrastructure (PKI) where each party $P_{i}$ possesses a long-term secret key $\textrm{sk}_{i}$ and a corresponding public key $\textrm{pk}_{i}$ for an IND-CPA secure public-key encryption scheme $(\texttt{Enc},\texttt{Dec})$ (Section III-D), used for secure share distribution. Communication occurs via a broadcast channel. The protocol also utilizes a Non-Interactive Zero-Knowledge (NIZK) proof system $(\textrm{Setup},\textrm{Prove},\textrm{Verify})$ (Section III-E) which is assumed to be complete, computationally sound, and zero-knowledge, operating with a common reference string crs.
The goal of FDKG is for any subset of participating parties $\mathbb{D}\subseteq\mathbb{P}$ to jointly generate a global public key $\textbf{E}\in\mathbb{G}$ such that the corresponding secret key $\textbf{d}\in\mathbb{Z}_{q}$ (where $\textbf{E}=G^{\textbf{d}}$) is implicitly shared among them. Specifically, each participant $P_{i}\in\mathbb{D}$ generates a partial key pair $(d_{i},E_{i})$ where $E_{i}=G^{d_{i}}$. The global key pair is formed by aggregation: $\textbf{d}=\sum_{i\in\mathbb{D}}d_{i}$ and $\textbf{E}=\prod_{i\in\mathbb{D}}E_{i}$. Each participant $P_{i}$ uses a local $(t,k)$-Shamir Secret Sharing (SSS) scheme (Section III-B) to distribute its partial secret $d_{i}$ among a self-selected guardian set $\mathcal{G}_{i}\subset\mathbb{P}\setminus\{P_{i}\}$ of size $k$.
@@ IV-B The FDKG Protocol ($\Pi_{FDKG}$)
The protocol proceeds in two rounds: Distribution and Reconstruction.
IV-B1 Round 1: Distribution Phase
@@
    Partial Key Generation: Sample partial secret $d_{i}\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,\mathbb{Z}_{q}$. Compute $E_{i}=G^{d_{i}}$.
@@
    • Sample PKE randomness: $(k_{i,j},r_{i,j})\>{\leftarrow\!\!\mbox{\tiny${\$}$\normalsize}}\,\mathbb{Z}_{q}\times\mathbb{Z}_{q}$.
    • Encrypt: $C_{i,j}\leftarrow\texttt{Enc}_{\textrm{pk}_{j}}(s_{i,j},[k_{i,j},r_{i,j}])$.
@@
    Proof Generation: Let the statement be $x_{i}=(E_{i},\{\textrm{pk}_{j}\}_{P_{j}\in\mathcal{G}_{i}},\mathbb{C}_{i})$. Let the witness be $w_{i}=(d_{i},\{s_{i,j}\}_{P_{j}\in\mathcal{G}_{i}},\mathbb{R}_{i})$. Generate $\pi_{FDKG_{i}}\leftarrow\textrm{Prove}(\textrm{crs},x_{i},w_{i})$.
@@ IV-B2 Round 2: Reconstruction Phase
    • Generate $\pi_{PS_{i}}\leftarrow\textrm{Prove}(\textrm{crs},E_{i},d_{i})$.
@@
    • Decrypt share: $s_{j,i}\leftarrow\texttt{Dec}(\textrm{sk}_{i},C_{j,i})$.
    • Generate proof: $\pi_{SPS_{j,i}}\leftarrow\textrm{Prove}(\textrm{crs},(C_{j,i},\textrm{pk}_{i},s_{j,i}),\textrm{sk}_{i})$.
@@
    Compute Global Key: If $D$ contains a valid entry for every $P_{i}\in\mathbb{D}$, compute $\textbf{d}=\sum_{d_{i}\in D}d_{i}\pmod{q}$. Output d. Else output failure.
@@ IV-C Communication Complexity
  • Round 1 (Distribution): Each participant $P_{i}\in\mathbb{D}$ broadcasts its partial public key $E_{i}$ (1 group element), $k$ encrypted shares $\{C_{i,j}\}$ (size depends on PKE scheme, in our scheme $k\times(2|\mathbb{G}|+|\mathbb{Z}_{q}|)$), and one NIZK proof $\pi_{FDKG_{i}}$. Assuming $|\mathbb{D}|\approx n$, the complexity is roughly $O(nk\cdot|C_{i,j}|+n\cdot|\pi_{FDKG}|)$. In our BabyJub/Groth16 instantiation, this leads to $O(nk)$ complexity dominated by encrypted shares.
@@
FDKG offers $O(n\cdot k)$ complexity for distribution, which is similar to standard DKG if $k\approx n$. Reconstruction complexity varies but can reach $O(n^{2})$ in worst-case guardian configurations.
@@ IV-D FDKG Example
Consider $n=10$ potential parties $\{P_{1},\dots,P_{10}\}$, local threshold $t=2$, and guardian set size $k=3$. Suppose $\mathbb{D}=\{P_{1},P_{3},P_{5},P_{7},P_{9}\}$ participate validly in Round 1.
@@
  • Round 1 (as per Section IV-B1): $P_{1}$ chooses $\mathcal{G}_{1}=\{P_{2},P_{3},P_{5}\}$, generates $(d_{1},E_{1})$, shares $d_{1}$ into $s_{1,2},s_{1,3},s_{1,5}$, encrypts them to get $\mathbb{C}_{1}=\{C_{1,2},C_{1,3},C_{1,5}\}$, generates $\pi_{FDKG_{1}}$, and broadcasts. Other participants in $\mathbb{D}$ do similarly. Observers compute $\textbf{E}=\prod_{P_{i}\in\mathbb{D}}E_{i}$.
  • Round 2 (Reconstruction, as per Section IV-B2): Suppose $\mathbb{T}=\{P_{1},P_{2},P_{3},P_{4},P_{5},P_{7}\}$.
    • $P_{1}$ broadcasts $(P_{1},\text{secret},d_{1},\pi_{PS_{1}})$.
    • $P_{2}$ (guardian for $P_{1}$) decrypts $C_{1,2}$ (using $\textrm{sk}_{2}$) to get $s_{1,2}$. Broadcasts $(P_{2},\text{share},P_{1},s_{1,2},\pi_{SPS_{1,2}})$.
    • (Similar actions for other talliers and shares).
  • Offline Reconstruction: An observer collects all valid broadcasted secrets and shares.
    • Recovers $d_{1}$ directly.
    • For $P_{5}$ (assuming $P_{5}\notin\mathbb{T}$ or didn’t reveal $d_{5}$): If $P_{3}$ (in $\mathcal{G}_{5}$) revealed $s_{5,3}$ and $P_{7}$ (in $\mathcal{G}_{5}$) revealed $s_{5,7}$ (both with valid proofs), the observer uses these two shares ($t=2$) to $\textrm{Reconstruct}(\{P_{3},P_{7}\},\{s_{5,3},s_{5,7}\})$ yielding $d_{5}^{\prime}$. Verifies $G^{d_{5}^{\prime}}=E_{5}$.
    • After recovering all $d_{i}$ for $P_{i}\in\mathbb{D}$, compute $\textbf{d}=\sum d_{i}$.

### v3

@@ III-A Notation
Let $\lambda$ denote the computational security parameter. We work within a cyclic group $\mathbb{G}$ of prime order $q$ generated by $G\in\mathbb{G}$. Group operations are written multiplicatively, and scalar multiplication is denoted by $G^{a}$ for $a\in\mathbb{Z}_{q}$. We assume the Discrete Logarithm Problem (DLP) is hard in $\mathbb{G}$. All algorithms are assumed to be Probabilistic Polynomial-Time (PPT) unless stated otherwise. Sampling a random element $x$ from a set $S$ is denoted by $x\>{\leftarrow\!\!\mbox{\tiny${\mathdollar}$\normalsize}}\,S$. We denote the set of potential parties as $\mathbb{P}=\{P_{1},\dots,P_{n}\}$.
@@ III-C Verifiable and Publicly Verifiable Secret Sharing (VSS/PVSS)
VSS schemes enhance SSS by allowing parties to verify the consistency of their shares relative to some public information. PVSS strengthens this by enabling any entity to perform this verification using only public data [35, 36]. PVSS schemes typically involve the dealer encrypting each share $s_{i}$ under the recipient $P_{i}$’s public key $\textrm{pk}_{i}$ and publishing these ciphertexts $C_{i}=\textrm{Enc}(\textrm{pk}_{i},s_{i})$ along with public commitments to the secret polynomial (e.g., $\{A_{k}=G^{a_{k}}\}$) and a public proof $\pi$ of consistency. Our FDKG protocol employs a PVSS scheme where each participant acts as a dealer.
@@ III-D Public Key Encryption (PKE)
  • $(\textrm{pk}_{i},\textrm{sk}_{i})\>{\leftarrow\!\!\mbox{\tiny${\mathdollar}$\normalsize}}\,\textrm{KeyGen}(1^{\lambda})$: Generates a public key $\textrm{pk}_{i}$ and a secret key $\textrm{sk}_{i}$.
@@ III-E Non-Interactive Zero-Knowledge (NIZK) Proofs
  • $\textrm{crs}\>{\leftarrow\!\!\mbox{\tiny${\mathdollar}$\normalsize}}\,\textrm{Setup}(1^{\lambda},\mathcal{R})$: Generates the public CRS.
  • $\pi\>{\leftarrow\!\!\mbox{\tiny${\mathdollar}$\normalsize}}\,\textrm{Prove}(\textrm{crs},x,w)$: Produces proof $\pi$.
@@ III-F Encrypted Channels for Share Distribution
Protocols like PVSS and DKG often require confidential transmission of secret shares ($s_{i}$) to specific recipients ($P_{i}$) over a public broadcast channel. This is achieved using an IND-CPA secure PKE scheme. Each potential recipient $P_{i}$ has a key pair $(\textrm{pk}_{i},\textrm{sk}_{i})$ where $\textrm{pk}_{i}$ is public. To send $s_{i}$ to $P_{i}$, the sender computes $c=\textrm{Enc}(\textrm{pk}_{i},s_{i},\textrm{rand})$ and broadcasts $c$. Only $P_{i}$ can compute $s_{i}=\textrm{Dec}(\textrm{sk}_{i},c)$.
@@
1Input: Recipient public key $\textrm{pk}_{i}$, scalar message $m\in\mathbb{Z}_{q}$, randomness $(k,r)\in\mathbb{Z}_{q}\times\mathbb{Z}_{q}$
@@ 5 $C_{2}=(\textrm{pk}_{i})^{k}\cdot M$
Algorithm 1 ElGamal Encryption $\textrm{Enc}(\textrm{pk}_{i},m,[k,r])$
1Input: Recipient secret key $\textrm{sk}_{i}$, ciphertext $(C_{1},C_{2},\Delta)$
@@ 3 $M=C_{2}\cdot(C_{1}^{\textrm{sk}_{i}})^{-1}$
Algorithm 2 ElGamal Decryption $\textrm{Dec}(\textrm{sk}_{i},(C_{1},C_{2},\Delta))$
@@ III-G Summary of Key Notations
Table II: Key Notations Used in the FDKG Protocol
@@
$\textrm{pk}_{i},\textrm{sk}_{i}$ PKE public/secret key pair for $P_{i}$.
@@
$\mathcal{G}_{i}$ Guardian set chosen by participant $P_{i}$.
$t$ Threshold for share reconstruction within a guardian set.
@@
$\mathbb{D}$ Set of valid parties in the generation phase.
@@
$\textrm{Enc}(\textrm{pk},\cdot,\cdot)$ PKE encryption algorithm.
$\textrm{Dec}(\textrm{sk},\cdot)$ PKE decryption algorithm.
@@
$B_{i},v_{i},r_{i}$ Voter $P_{i}$’s encrypted ballot, encoded vote,
and blinding factor.
@@
Our implementation instantiates IND-CPA PKE with an ElGamal variant over BabyJub (Edwards curve) and NIZKs with Groth16. The DLP hardness assumption is therefore the discrete-log assumption in the BabyJub subgroup; IND-CPA security reduces to the DDH assumption in that group. Groth16 soundness/zero-knowledge rely on pairing-friendly curve security (BN254 in our implementation) and a trusted SRS for the given circuits. The security reductions in Section V use only these abstract properties; other IND-CPA PKEs or NIZKs (PLONK/Marlin/STARKs) could be substituted with corresponding trade-offs in proof size and proving time.
@@ IV Federated Distributed Key Generation (FDKG)
Federated Distributed Key Generation (FDKG) generalizes DKG to dynamic participation: any subset of parties can generate a joint key and delegate reconstruction to self-chosen guardian sets.
@@ IV-A Protocol Setting and Goal
We consider a set $\mathbb{P}=\{P_{1},\dots,P_{n}\}$ of $n$ potential parties. The system operates over a cyclic group $\mathbb{G}$ of prime order $q$ with generator $G$. We assume a Public Key Infrastructure (PKI) where each party $P_{i}$ possesses a long-term secret key $\textrm{sk}_{i}$ and a corresponding public key $\textrm{pk}_{i}$ for an IND-CPA secure public-key encryption scheme $(\textrm{Enc},\textrm{Dec})$ (Section III-D), used for secure share distribution. Communication occurs via a broadcast channel. The protocol also utilizes a Non-Interactive Zero-Knowledge (NIZK) proof system $(\textrm{Setup},\textrm{Prove},\textrm{Verify})$ (Section III-E) which is assumed to be complete, computationally sound, and zero-knowledge, operating with a common reference string crs.
The goal of FDKG is for any subset of participating parties $\mathbb{D}\subseteq\mathbb{P}$ to jointly generate a global public key $\textbf{E}\in\mathbb{G}$ such that the corresponding secret key $\textbf{d}\in\mathbb{Z}_{q}$ (where $\textbf{E}=G^{\textbf{d}}$) is shared among them. Specifically, each participant $P_{i}\in\mathbb{D}$ generates a partial key pair $(d_{i},E_{i})$ where $E_{i}=G^{d_{i}}$. The global key pair is formed by aggregation: $\textbf{d}=\sum_{i\in\mathbb{D}}d_{i}$ and $\textbf{E}=\prod_{i\in\mathbb{D}}E_{i}$. Each participant $P_{i}$ uses a local $(t,k)$-Shamir Secret Sharing (SSS) scheme (Section III-B) to distribute its partial secret $d_{i}$ among a self-selected guardian set $\mathcal{G}_{i}\subset\mathbb{P}\setminus\{P_{i}\}$ of size $k$.
@@ IV-B The FDKG Protocol ($\Pi_{FDKG}$)
The protocol proceeds in two rounds: Generation and Reconstruction.
IV-B1 Round 1: Generation phase
@@
    Partial Key Generation: Sample partial secret $d_{i}\>{\leftarrow\!\!\mbox{\tiny${\mathdollar}$\normalsize}}\,\mathbb{Z}_{q}$. Compute $E_{i}=G^{d_{i}}$.
@@
    • Sample PKE randomness: $(k_{i,j},r_{i,j})\>{\leftarrow\!\!\mbox{\tiny${\mathdollar}$\normalsize}}\,\mathbb{Z}_{q}\times\mathbb{Z}_{q}$.
    • Encrypt: $C_{i,j}\leftarrow\textrm{Enc}_{\textrm{pk}_{j}}(s_{i,j},[k_{i,j},r_{i,j}])$.
@@
    Proof Generation: Let the public statement be $x_{i}=(E_{i},\{\textrm{pk}_{j}\}_{P_{j}\in\mathcal{G}_{i}},\mathbb{C}_{i})$. Let the witness be $w_{i}=(d_{i},\{s_{i,j}\}_{P_{j}\in\mathcal{G}_{i}},\mathbb{R}_{i})$. Generate $\pi_{FDKG_{i}}\leftarrow\textrm{Prove}(\textrm{crs},x_{i},w_{i})$ (see Appendix B-A).
@@ IV-B2 Round 2: Reconstruction Phase
    • Generate $\pi_{PS_{i}}\leftarrow\textrm{Prove}(\textrm{crs},E_{i},d_{i})$ (see Appendix B-B).
@@
    • Decrypt share: $s_{j,i}\leftarrow\textrm{Dec}(\textrm{sk}_{i},C_{j,i})$.
    • Generate proof: $\pi_{SPS_{j,i}}\leftarrow\textrm{Prove}(\textrm{crs},(C_{j,i},\textrm{pk}_{i},s_{j,i}),\textrm{sk}_{i})$ (see Appendix B-C).
@@
    Compute Global Key: If $D$ contains a valid entry for every $P_{i}\in\mathbb{D}$, compute $\textbf{d}=\sum_{d_{i}\in D}d_{i}$. Output d. Else output failure.
@@ IV-C Communication Complexity
  • Round 1 (Generation): Each participant $P_{i}\in\mathbb{D}$ broadcasts its partial public key $E_{i}$ (1 group element), $k$ encrypted shares $\{C_{i,j}\}$ (size depends on PKE scheme, in our scheme $k\times(2|\mathbb{G}|+|\mathbb{Z}_{q}|)$), and one NIZK proof $\pi_{FDKG_{i}}$. Assuming $|\mathbb{D}|\approx n$, the complexity is roughly $O(nk\cdot|C_{i,j}|+n\cdot|\pi_{FDKG}|)$. In our BabyJub/Groth16 instantiation, this leads to $O(nk)$ complexity dominated by encrypted shares.
@@
FDKG offers $O(n\cdot k)$ complexity for distribution, which is similar to standard DKG if $k\approx n$. Reconstruction complexity varies but can reach $O(n^{2})$ in worst-case guardian—when a single party is selected as a guardian by every other participant and must reveal all n1n-1 shares.
@@ IV-D FDKG Example
Consider $n=10$ potential parties $\{P_{1},\dots,P_{10}\}$, local threshold $t=2$, and guardian set size $k=3$. Suppose $\mathbb{D}=\{P_{1},P_{3},P_{5},P_{7},P_{9}\}$ participate correctly in Round 1.
@@
  • Round 1 Online
    • $P_{1}$ chooses $\mathcal{G}_{1}=\{P_{2},P_{3},P_{5}\}$.
    • Generates $(d_{1},E_{1})$.
    • Shares $d_{1}$ into $s_{1,2},s_{1,3},s_{1,5}$.
    • Encrypts shares to get $\mathbb{C}_{1}=\{C_{1,2},C_{1,3},C_{1,5}\}$.
    • Generates $\pi_{FDKG_{1}}$ and broadcasts. Others in $\mathbb{D}$ do similarly.
  • Round 1 Offline
    • Once broadcasts are received, everyone computes $\textbf{E}=\prod_{P_{i}\in\mathbb{D}}E_{i}$.
  • Round 2 Online
    • Suppose $\mathbb{T}=\{P_{3},P_{5},P_{7}\}$.
    • $P_{3},P_{5},P_{7}$ each broadcast their $(\text{secret},d_{\cdot},\pi_{PS_{\cdot}})$.
    • As guardians, they also decrypt and broadcast the necessary shares for offline parties:
        $P_{3}$ and $P_{5}$ (guardians for $P_{1}$) broadcast $(\text{share},P_{1},s_{1,3},\pi_{SPS_{1,3}})$ and $(\text{share},P_{1},s_{1,5},\pi_{SPS_{1,5}})$ respectively.
        $P_{5}$ and $P_{7}$ (guardians for $P_{9}$) broadcast $(\text{share},P_{9},s_{9,5},\pi_{SPS_{9,5}})$ and $(\text{share},P_{9},s_{9,7},\pi_{SPS_{9,7}})$ respectively.
  • Round 2 Offline
    • Collect all valid broadcasts: direct secrets from $P_{3},P_{5},P_{7}$ and guardian shares for $P_{1}$ and $P_{9}$.
    • Reconstruct offline parties using $t=2$ shares:
        For $P_{1}$: use $\{(P_{3},s_{1,3}),(P_{5},s_{1,5})\}$ to compute $d_{1}^{\prime}\leftarrow\textrm{Reconstruct}(\{P_{3},P_{5}\},\{s_{1,3},s_{1,5}\})$ and verify $G^{d_{1}^{\prime}}=E_{1}$.
        For $P_{9}$: use $\{(P_{5},s_{9,5}),(P_{7},s_{9,7})\}$ to compute $d_{9}^{\prime}\leftarrow\textrm{Reconstruct}(\{P_{5},P_{7}\},\{s_{9,5},s_{9,7}\})$ and verify $G^{d_{9}^{\prime}}=E_{9}$.
    • With $d_{3},d_{5},d_{7}$ revealed directly and $d_{1}^{\prime},d_{9}^{\prime}$ reconstructed, compute $\textbf{d}=\sum_{P_{i}\in\mathbb{D}}d_{i}$.